OpenStudy (anonymous):

What is the slope intercept form of the equation x - 3y = 3?

OpenStudy (anonymous):

A. y = -3x + 2 B. y = -1/3x - 2 C. y = 3x + 6 D. y = 1/3x - 1

OpenStudy (uri):

same like before @fukaawa

OpenStudy (uri):

y=mx+b arrange the equation

OpenStudy (uri):

-3y=x+3

OpenStudy (uri):

y=1/-3x+3/-3